# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

### v1

@@ 1.5. Scope of this Survey
Several existing reviews have been reported on HPE and APE separately. Prior surveys (Zheng et al., 2020; Ji and Liu, 2009; Sarafianos et al., 2016; Khan and Wan, 2018; Gong et al., 2016; Liu et al., 2022a; Song et al., 2021; Dubey and Dixit, 2023; Wang et al., 2021b; Moeslund and Granum, 2001; Moeslund et al., 2006; Poppe, 2007) have primarily concentrated on vision-based human motion capture, encompassing aspects of pose estimation, motion tracking, and action recognition. While these studies have significantly contributed to HPE, there remains a conspicuous absence of in-depth comparison and analysis between HPE and APE techniques. Reference (Jiang et al., 2022) discusses vision-based APE between 2013 and 2021, yet it falls short of offering an exhaustive discussion on the existing challenges and future directions in APE. Another limitation is that only a limited number of reviews (Zheng et al., 2020) discuss HPE or APE from sources other than vision-based inputs. This paper aims to bridge these gaps in the literature by offering a detailed analysis that extends beyond the scope of prior surveys and reviews:
  1. (1)
    We meticulously reviewed 178 papers up to 2024, categorizing them based on different parameters such as their output forms (e.g. 2D vs. 3D pose estimation; keypoint vs. mesh pose representation), input sensor and modality types (monocular inputs vs. other sensory inputs; single-modal APE vs. multi-modal APE), learning paradigms (e.g. heatmap-based methods vs. regression methods; top-down methods vs. bottom-up methods; supervised vs. semi-supervised vs. self-supervised vs. unsupervised APE methods), experimental setup (e.g. single-camera-view vs. multiple-camera-views approaches; indoor vs. in-the-wild setup; the range of the sensors).
  2. (2)
@@
Table 2. RGB-based APE Dataset (Images-based). Column N indicates no. of animals: ”M” denotes multiple animals’ pose is annotated in one frame, while ”S” denotes a single animal’s pose is annotated. Column I/W is ”indoor” [Uncaptioned image] or ”in-the-wild” [Uncaptioned image] set-up.
Sl Year Datasets & Size Animal Classes Source/Experiment N I/W Annotation
set-up
@@
1 2011 Stanford Dogs (Khosla et al., 2011) Dogs (120 breeds) a. ImageNet data N/A [Uncaptioned image] Class labels,
20k images b. Single view bounding box
2 2014 COCO (Lin et al., 2014b) Cat, Dog, Horse, Bear, a. Online Repositories N/A [Uncaptioned image] & [Uncaptioned image] Bounding box, class
1k images for animals Sheep, Zebra, Giraffe b. Single view label, segmentation
3 2018 Animals-10 (C., 2018) Dog, Cat, Horse, a. Google Images N/A [Uncaptioned image] & [Uncaptioned image] N/A
28k images Sheep, Cow, Elephant b. Single view
4 2019 Animal Pose (Cao et al., 2019) Dogs, Cats, Cows, a. Animals-10 (C., 2018) & S [Uncaptioned image] & [Uncaptioned image] 20 keypoints (2D pose),
4k images Horses, Sheep, COCO (Lin et al., 2014b) bounding box
& other 7 categories b. Single view
5 2019 Grevy’s Zebra (Zuffi et al., 2019b) Synthetic Zebra a. Computer generated S [Uncaptioned image] 28 keypoints (2D pose),
@@
6 2020 Synthetic Animal (Mu et al., 2020) Synthetic Hound, a. COCO2017 (Lin et al., 2014b) S [Uncaptioned image] & [Uncaptioned image] 18 keypoints (2D &
10k images Horse, Sheep, b. Single view 3D pose), depth maps,
Elephant, Tiger segmentation
7 2020 StanfordExtra (Biggs et al., 2020) Dogs a. Stanford Dogs (Khosla et al., 2011) S [Uncaptioned image] 24 keypoints (2D pose),
@@
8 2021 AP-10K (Yu et al., 2021b) 23 animal families, a. Online Repositories [Uncaptioned image] 17 keypoints (2D pose)
10,015 images 54 species b. Single view M
9 2021 MacaquePose(Labuguen et al., 2021) Macaque Monkey a. Manually collected M [Uncaptioned image] 17 keypoint (2D pose)
13,000 images & Onlince
@@
10 2022 Animal Kingdom (Ng et al., 2022b) 850 animal species a. Online Repositories S [Uncaptioned image] 23 keypoints (2D pose)
@@
11 2022 OpenMonkey 26 Monkey species a. Manually collected S [Uncaptioned image] 17 keypoints (2D pose)
Challenge (Yuan Yao and Park, 2022) & Online
@@
12 2023 OpenApePose (Desai et al., 2023) 6 Ape species a. Manually collected S [Uncaptioned image] 16 keypoints (2D pose)
@@
13 2023 Animal3D (Xu et al., 2024) 40 Mammal species a. PartImageNet (He et al., 2022) S [Uncaptioned image] 26 keypoints (2D pose),
3379 images &   COCO (Lin et al., 2014b) 3D mesh
@@ 2. Datasets
2.1. RGB-based Animal Pose Estimation Dataset
@@
2.1.1. Images-based Animal Pose Estimation Dataset
Early datasets in this category lacked large-scale annotated animal pose labels. For example, the Stanford Dogs (2011) dataset (Khosla et al., 2011) includes 20k images of 120 dog breeds with class labels and bounding boxes but no pose keypoints. Similarly, the COCO (2014) dataset (Lin et al., 2014b) with 1k images only provides annotations such as bounding boxes, segmentation, and class labels for animals like cats, dogs, and horses. The Animals-10 (2018) dataset (C., 2018), consisting of 14k images of common animals such as dogs, cats, and horses, also lacks animal keypoint annotations.
As there are few pose-labeled animal datasets available, more 2D pose-labeled animal datasets have been developed to evaluate the performance of APE objectively. For instance, the Animal Pose (2019) (Cao et al., 2019) consists of 4k images across 10 animal categories, with each image annotated with 20 2D keypoints to represent the animal skeleton. The StanfordExtra (2020) (Biggs et al., 2020) offers 12k dogs images with 24 2D pose keypoints. The AP-10K (2021) (Yu et al., 2021b) provides 10k images from 23 animal families, with 17 2D keypoints labeled for each detected animal. The recent Animal Kingdom (2022) (Ng et al., 2022b), the OpenMonkeyChallenge (2022) (Yuan Yao and Park, 2022), and the OpenApePose (2023) (Desai et al., 2023) all include annotated 2D animal poses keypoints.
Beyond 2D keypoint pose labels, more recent datasets have started providing 3D pose annotations, including 3D keypoints and meshes. Examples include the Synthetic Grevy’s Zebra (2019) (Zuffi et al., 2019b), the Synthetic Animal (2020) (Mu et al., 2020), and the Animal3D (2023) (Xu et al., 2024), which offer both high-quality 2D keypoint annotations and 3D pose labels. This development reflects the growing diversity and depth of datasets available for APE research.
2.1.2. Video-based Animal Pose Estimation Dataset
Table 3. RGB-based APE Dataset (Video-based). Column [Uncaptioned image] Rng is camera range. NGD denotes National Geographic Documentaries; the symbol [Uncaptioned image] denotes cameras. All the other symbols and column meanings are defined in table 2.
Sl Year Datasets & Size Animal Classes Source & Experiment set-up [Uncaptioned image] Rng N I/W Annotation
@@ 2.1.2. Video-based Animal Pose Estimation Dataset
14 2016 TigDog (Del Pero et al., 2017, 2015) Dogs, Horses, a. YouTube & NGD videos N/A S [Uncaptioned image] 19 keypoints
110k frames Tigers b. Single view (2D pose)
55 videos
15 2018 BADJA (Biggs et al., 2018) Bears, Camels, a. DAVIS (Pont-Tuset et al., 2017) videos 1-20m S [Uncaptioned image] 20 keypoints
11 videos Dogs, Horses, b. Single view (2D & 3D pose),
Tigers, Cats, Cows segmentation
16 2020 ATRW (Li et al., 2020b) Amur tigers a. Manually collected N/A M [Uncaptioned image] 15 keypoints
@@
(Bala et al., 2020) 195,228 frames b. Multi views (62 [Uncaptioned image] views) (3D pose)
18 2021 AcinoSet (Joska et al., 2021) Cheetahs a. Manually collected 3-12m S [Uncaptioned image] 20 keypoints
119k frames b. Multi views (2D & 3D pose)
93 videos (6 GoPro [Uncaptioned image] viewpoint)
19 2021 Horse-30 (Mathis et al., 2021) Horses a. Manually collected N/A S [Uncaptioned image] 22 keypoints
@@
20 2022 APT-36K (Yang et al., 2022b) Quadruped a. YouTube videos N/A M [Uncaptioned image] 17 keypoints
36k frames Mammals b. Single view (2D pose)
2.4k videos (30 Species)
21 2023 ChimpACT (Ma et al., 2023) Chimpanzees a. Manually collected N/A M [Uncaptioned image] & [Uncaptioned image] 16 keypoints
16,028 frames b. Single view (2D pose)
163 videos
@@
We have listed several well-known datasets in this category, with most pose annotations being presented in 2D and 3D keypoint formats (table 3). In particular, most 2D keypoint pose labels in these datasets (e.g. Horse-30 (Mathis et al., 2021), APT-36K (Yang et al., 2022b), ChimpACT (Ma et al., 2023)) are provided by well-trained annotators. For instance, the TigDog (2016) (Del Pero et al., 2017, 2015) includes 110k frames of in-the-wild dogs, horses, and tigers, with 19 keypoint annotations for 2D poses. The Horse-30 (2021) (Mathis et al., 2021) provides 8k frames of horses in-the-wild, with 22 annotated 2D keypoints. While these datasets are valuable, they are limited to individual animals. However, there is a recent growing trend toward providing multi-animal pose annotations (e.g. ATRW (Li et al., 2020b), APT-36K (Yang et al., 2022b) and ChimpACT (Ma et al., 2023)) to account for the frequent interactions between quadruped animals.
@@
There is also a small collection of publicly available 3D animal pose benchmark datasets. For example, the BADJA (2018) (Biggs et al., 2018) contains 11 in-the-wild animal videos annotated with 20 3D joints mapped onto a SMAL mesh. The AcinoSet (2021) (Joska et al., 2021) offers 7k frames of hand-labelled 3D keypoints from cheetahs, capturing their running behaviour in-the-wild. Additionally, the OpenMonkeyPose (2020) (Bala et al., 2020) includes 195k well-annotated 3D keypoint pose instances, spanning a large variation of poses and positions of monkeys across diverse activities. It is worth noting that both the OpenMonkeyPose (Bala et al., 2020) and AcinoSet (Joska et al., 2021) include synchronized multiple-camera-views footage and camera calibration, enabling multiple-camera-views geometry and triangulation for precise 3D pose estimation (e.g. OpenMonkeyPose offers 62-viewpoint, while AcinoSet uses 6). We also discuss using MoCap systems with markers to obtain ground truth 3D animal poses in the following subsection.
@@
One common limitation across many datasets in this category is their focus on specific animal species. Due to the significant variations in appearance, behaviour, and joint distributions between species, APE models trained on these datasets often fail to generalize to unseen species, leading to poor performance in real-world scenarios. To fill this gap, the APT-36K (2022) (Yang et al., 2022b) was introduced, covering 2k video clips of 30 quadruped species with high-quality 2D keypoint annotations. This enables inter-species generalization testing and reflects the need for models that work across diverse species.
@@
Table 4. Other-Sensor-based APE Dataset. [Uncaptioned image] is markers; vol. is volume. All the other symbols and column meanings are defined in table 2 and table 3.
@@
22 2021 PAIR-R24M (Marshall et al., 2021) Rat a. Manually collected N/A M [Uncaptioned image] 12 keypoints
@@
23 2021 Rat7M (Dunn et al., 2021) Rat a. Manually collected N/A S [Uncaptioned image] 20 keypoints
@@
24 2023 3D-POP (Naik et al., 2023) Pigeons a. Manually collected 0-4.2m (RGB) M [Uncaptioned image] 9 keypoints
300k frames b. Multi views (4 RGB [Uncaptioned image]) 15x7x4m vol. (2D & 3D pose)
4 videos c. 30 Mocap [Uncaptioned image] & 8 [Uncaptioned image] (Mocap)
25 2024 PFERD (Li et al., 2024c) Horses a. Manually collected 16x20m S [Uncaptioned image] 3D Mesh,
@@
26 2020 RGBD-Dog (Kearney et al., 2020a) Dogs a. Manually collected 0.5-4.5m S [Uncaptioned image] 3D Mesh,
@@
27 2024 WildPose (Muramatsu et al., 2024) Lion, a. Manually collected 2–120m M [Uncaptioned image] 15 keypoints,
@@
28 2023 LoTE-Animal (Liu et al., 2023) 11 animal a. Manually collected & Online N/A S [Uncaptioned image] 17 keypoints
@@
29 2024 VAREN (Zuffi et al., 2024) Horses a. Manually collected N/A S [Uncaptioned image] 3D Mesh
@@
30 2024 Yigit et al. (Yigit et al., 2022) Horse a. Manually collected S [Uncaptioned image] & [Uncaptioned image] 3D Motion data
@@
2.2. Other-Sensor-based Animal Pose Estimation Dataset

### v2

@@ 1.5 Scope of this Survey
Several existing reviews have been reported on HPE and APE separately. Prior surveys [171, 51, 125, 57, 38, 80, 131, 29, 143, 93, 94, 115, 173] have primarily concentrated on vision-based human motion capture, encompassing aspects of pose estimation, motion tracking, action recognition and motion generation. While these studies have significantly contributed to HPE, there remains a conspicuous absence of in-depth comparison and analysis between HPE and APE techniques. Reference [52] discusses vision-based APE between 2013 and 2021, yet it falls short of offering an exhaustive discussion on the existing challenges and future directions in APE. Another limitation is that only a limited number of reviews [171] discuss HPE or APE from sources other than vision-based inputs. This paper aims to bridge these gaps in the literature by offering a detailed analysis that extends beyond the scope of prior surveys and reviews:
  • We meticulously reviewed 176 papers up to 2024, categorizing them based on different parameters such as their output forms (e.g. 2D vs. 3D pose estimation; keypoint vs. mesh pose representation), input sensor and modality types (monocular inputs vs. other sensory inputs; uni-modal APE vs. multi-modal APE), learning paradigms (e.g. heatmap-based methods vs. regression methods; top-down methods vs. bottom-up methods; supervised vs. semi-supervised vs. self-supervised vs. unsupervised APE methods), and experimental setup (e.g. single-camera-view vs. multiple-camera-views approaches; indoor vs. in-the-wild setup; the range of the sensors).
@@
2 Datasets
Table 2: RGB-based APE Dataset (Images-based). Column N indicates no. of animals: M denotes multiple animals’ pose is annotated in one frame, while S denotes a single animal’s pose is annotated. In I/W column, [Uncaptioned image] and [Uncaptioned image] represent indoor and in-the-wild setups, respectively. Exp. is Experiment. The dataset citation is available in column Year.
Sl Year Datasets & Size Animal Classes Source & Exp. Set-up N I/W Annotation
@@ 2 Datasets
1 2011 Stanford Dogs Dogs (120 breeds) a. ImageNet data N/A [Uncaptioned image] Class label, bounding box
20k images b. Single view
2 2014 COCO Bear, Sheep, Zebra, a. Online repositories N/A [Uncaptioned image] & [Uncaptioned image] Bounding box, class label,
1k animals images Cat, Dog, Horse, b. Single view segmentation
Giraffe
3 2018 Animals-10 Dog, Horse, Sheep, a. Google images N/A [Uncaptioned image] & [Uncaptioned image] N/A
28k images Cat, Cow, Elephant b. Single view
4 2019 Animal Pose Dogs, Cats, Cows, a. Animals-10 (2018) & S [Uncaptioned image] & [Uncaptioned image] 20 keypoints (2D pose),
4k images Horses, Sheep & COCO (2014) bounding box
other 7 categories b. Single view
5 2019 Grevy’s Zebra Synthetic Zebra a. Computer generated S [Uncaptioned image] 28 keypoints (2D pose),
@@
6 2020 Synthetic Animal Synthetic Hound, a. COCO2017 (2014) S [Uncaptioned image] & [Uncaptioned image] 18 keypoints (2D&3D pose),
10k images Elephant, Tiger, b. Single view depth maps, segmentation
Horse, Sheep
7 2021b AP-10K 23 animal families, a. Online repositories M [Uncaptioned image] 17 keypoints (2D pose)
10,015 images 54 species b. Single view
8 2020 StanfordExtra Dogs a. Stanford Dogs (2011) S [Uncaptioned image] 24 keypoints (2D pose),
@@
9 2021 MacaquePose Macaque Monkey a. Manually collected & M [Uncaptioned image] 17 keypoint (2D pose)
13,000 images Online
@@
10 2022 Animal Kingdom 850 animal species a. Online repositories S [Uncaptioned image] 23 keypoints (2D pose)
@@
11 2022 OpenMonkey 26 Monkey species a. Manually collected & S [Uncaptioned image] 17 keypoints (2D pose)
Challenge Online
@@
12 2023 OpenApePose 6 Ape species a. Manually collected S [Uncaptioned image] 16 keypoints (2D pose)
@@
13 2023 Animal3D 40 Mammal species a. PartImageNet (2022) & S [Uncaptioned image] 26 keypoints (2D pose),
3379 images COCO (2014) 3D mesh
@@
2.1 RGB-based Animal Pose Estimation Dataset
@@
2.1.1 Images-based Animal Pose Estimation Dataset
Early datasets in this category lacked large-scale annotated animal pose labels. For example, the Stanford Dogs (2011) dataset [58] includes 20k images of 120 dog breeds with class labels and bounding boxes but no pose keypoints. Similarly, the COCO (2014) dataset [75] with 1k images only provides annotations such as bounding boxes, segmentation, and class labels for animals like cats, dogs, and horses. The Animals-10 (2018) dataset [1], consisting of 14k images of common animals such as dogs, cats, and horses, also lacks animal keypoint annotations.
As there are few pose-labeled animal datasets available, more 2D pose-labeled animal datasets have been developed to evaluate the performance of APE objectively. For instance, the Animal Pose (2019) [15] consists of 4k images across 10 animal categories, with each image annotated with 20 2D keypoints to represent the animal skeleton. The StanfordExtra (2020) [10] offers 12k dogs images with 24 2D pose keypoints. The AP-10K (2021) [164] provides 10k images from 23 animal families, with 17 2D keypoints labeled for each detected animal. The recent Animal Kingdom (2022) [103], the OpenMonkeyChallenge (2022) [158], and the OpenApePose (2023) [25] all include annotated 2D animal poses keypoints.
Beyond 2D keypoint pose labels, more recent datasets have started providing 3D pose annotations, including 3D keypoints and meshes. Examples include the Synthetic Grevy’s Zebra (2019) [176], the Synthetic Animal (2020) [95], and the Animal3D (2023) [150], which offer both high-quality 2D keypoint annotations and 3D pose labels. This development reflects the growing diversity and depth of datasets available for APE research.
2.1.2 Video-based Animal Pose Estimation Dataset
Table 3: RGB-based APE Dataset (Video-based). NGD denotes National Geographic Documentaries; the symbol [Uncaptioned image] denotes cameras. Column [Uncaptioned image] Rng is camera range. Exp. is Experiment. The dataset citation is available in column Year. All the other symbols and column meanings are defined in Table 2.
Sl Year Datasets & Size Animal Classes Source & Exp. Set-up [Uncaptioned image] Rng N I/W Annotation
@@ 2.1.2 Video-based Animal Pose Estimation Dataset
14 2015 TigDog Dogs, Horses, Tigers a. YouTube & NGD videos N/A S [Uncaptioned image] 19 keypoints
110k frames, 55 videos b. Single view (2D pose)
15 2018 BADJA Bears, Camels, Dogs, a. DAVIS (2017) videos 1-20m S [Uncaptioned image] 20 keypoints
11 videos Horses, Tigers, Cats, b. Single view (2D&3D pose),
Cows segmentation
16 2020 ATRW Amur tigers a. Manually collected N/A M [Uncaptioned image] 15 keypoints
@@
195,228 frames b. Multi views (62 [Uncaptioned image] views) (3D pose)
18 2021 AcinoSet Cheetahs a. Manually collected 3-12m S [Uncaptioned image] 20 keypoints
119k frames, 93 videos b. Multi views (6 [Uncaptioned image] views) (2D&3D pose)
19 2021 Horse-30 Horses a. Manually collected N/A S [Uncaptioned image] 22 keypoints
@@
20 2022 APT-36K Quadruped Mammals a. YouTube videos N/A M [Uncaptioned image] 17 keypoints
36k frames, 2.4k videos (30 Species) b. Single view (2D pose)
21 2023 ChimpACT Chimpanzees a. Manually collected N/A M [Uncaptioned image] & [Uncaptioned image] 16 keypoints
16,028 frames, 163 videos b. Single view (2D pose)
@@
We have listed several well-known datasets in this category, with most pose annotations being presented in 2D and 3D keypoint formats (Table 3). In particular, most 2D keypoint pose labels in these datasets (e.g. Horse-30 [91], APT-36K [157], ChimpACT [87]) are provided by well-trained annotators. For instance, the TigDog (2016) [23, 22] includes 110k frames of in-the-wild dogs, horses, and tigers, with 19 keypoint annotations for 2D poses. The Horse-30 (2021) [91] provides 8k frames of horses in-the-wild, with 22 annotated 2D keypoints. While these datasets are valuable, they are limited to individual animals. However, there is a recent growing trend toward providing multi-animal pose annotations (e.g. ATRW [73], APT-36K [157] and ChimpACT [87]) to account for the frequent interactions between quadruped animals.
@@
There is also a small collection of publicly available 3D animal pose benchmark datasets. For example, the BADJA (2018) [11] contains 11 in-the-wild animal videos annotated with 20 3D joints mapped onto a SMAL mesh. The AcinoSet (2021) [53] offers 7k frames of hand-labelled 3D keypoints from cheetahs, capturing their running behaviour in-the-wild. Additionally, the OpenMonkeyPose (2020) [8] includes 195k well-annotated 3D keypoint pose instances, spanning a large variation of poses and positions of monkeys across diverse activities. It is worth noting that both the OpenMonkeyPose [8] and AcinoSet [53] include synchronized multiple-camera-views footage and camera calibration, enabling multiple-camera-views geometry and triangulation for precise 3D pose estimation (e.g. OpenMonkeyPose offers 62-viewpoint, while AcinoSet uses 6). We also discuss using MoCap systems with markers to obtain ground truth 3D animal poses in the following subsection.
@@
One common limitation across many datasets in this category is their focus on specific animal species. Due to the significant variations in appearance, behaviour, and joint distributions between species, APE models trained on these datasets often fail to generalize to unseen species, leading to poor performance in real-world scenarios. To fill this gap, the APT-36K (2022) [157] was introduced, covering 2k video clips of 30 quadruped species with high-quality 2D keypoint annotations. This enables inter-species generalization testing and reflects the need for models that work across diverse species.
@@
Table 4: Other-Sensor-based APE Dataset. [Uncaptioned image] is markers; vol. is volume. The dataset citation is available in column Year. All the other symbols and column meanings are defined in Table 2 and Table 3.
@@
22 2021 PAIR-R24M Rat a. Manually collected N/A M [Uncaptioned image] 12 keypoints
@@
23 2021 Rat7M Rat a. Manually collected N/A S [Uncaptioned image] 20 keypoints
@@
24 2023 3D-POP Pigeons a. Manually collected M [Uncaptioned image] 9 keypoints
300k frames, b. Multi views (4 RGB [Uncaptioned image]) 0-4.2m (RGB) (2D & 3D pose)
4 videos c. 30 MoCap [Uncaptioned image] & 8 [Uncaptioned image] 15x7x4m (MoCap)
25 2024a PFERD Horses a. Manually collected 16x20m S [Uncaptioned image] 3D Mesh,
@@
26 2020 RGBD-Dog Dogs a. Manually collected 0.5-4.5m S [Uncaptioned image] 3D Mesh,
@@
27 2024 WildPose Lion, a. Manually collected 2–120m M [Uncaptioned image] 15 keypoints,
@@
28 2023 LoTE-Animal 11 animal a. Manually collected & Online N/A S [Uncaptioned image] 17 keypoints
@@
29 2024 VAREN Horses a. Manually collected N/A S [Uncaptioned image] 3D Mesh
@@
30 2022 Yigit et al. Horse a. Manually collected S [Uncaptioned image] & [Uncaptioned image] 3D Motion data
@@
2.2 Other-Sensor-based Animal Pose Estimation Dataset